8. A 0.0450-kg golf ball initially at rest is given a speed of 25.0 m/s when

it is struck by a club. If the club and ball are in contact for 2.00 ms, what

average force acts on the ball? Is the effect of the ball’s weight during

the time of contact significant? Why or why not?


1
Expert's answer
2021-10-10T16:00:28-0400

Given:

"m=0.0450\:\rm kg"

"v_i=0"

"v_f=25.0\:\rm m/s"

"\Delta t=0.00200\:\rm m/s"


The average force

"F=m\frac{\Delta v}{\Delta t}=0.0450*\frac{25.0}{0.00200}=563\:\rm N"

The weight of the ball

"W=mg=0.0450*9.81=0.44\:\rm N"

Since "W\ll F", so the weight isn't significant.